Each ball has a mass of 2.5kg & you can throw each at a velocity of 10m/s. The board has a mass of 20kg. The first is a tennis ball & will bonce back at a velocity of 8m/s. The second ball is clay & will stick to the board. Show conservation of momentum
1
Expert's answer
2012-11-08T08:42:27-0500
For the tennis ball: 2.5 kg *10 m/s = 2.5 kg * (-8) m/s + 20 kg * u,
where u = 2.5 kg * 18 m/s / 20 kg = 2.25 m/s is the eventual velocity of the board.

For the clay ball: 2.5 kg * 10 m/s = (2.5 kg + 20 kg) * u,
where u = 2.5 kg * 10 m/s / (2.5 kg + 20 kg) = 1.11 m/s is the eventual velocity of the board with the ball stuck.